OpenStudy (anonymous):

Use the discriminant to determine the nature of the roots of 2x2 = 0.

OpenStudy (saranya):

The discriminant is zero, so we should expect 1 real solution :)

OpenStudy (callisto):

discriminant = b^2 - 4ac = 0^2 - 4(2)(0) = 0 discriminant > 0 => 2 distinct real roots discriminant = 0 => a double real root discriminant < 0 => 2 distinct imaginary roots

OpenStudy (saranya):

Use the discriminant to find out the nature and number of solutions :)

OpenStudy (anonymous):

\[2x^2 = 0\] Here, a = 2, b= 0 and c = 0 As, \[D = \sqrt{b^2 - 4ac}\] D = 0.. So, this equation has equal and real roots..

OpenStudy (unklerhaukus):